OpenStudy (anonymous):

pls...find the value of A (trignometry) if; 2cos^2(A/2)=1

OpenStudy (anonymous):

cos x =0 after that you can find all answers by trig equations

OpenStudy (akashdeepdeb):

cos^2 (A/2) = (1 + cosA)/2 [An Identity] {Tell me if you want the proof} So \[\frac{ 1+cosA }{ 2 } = \frac{ 1 }{ 2 }\] cos A = 0 A = 90' Understood? :)

OpenStudy (anonymous):

\[2\cos^2(\frac{A}{2})=1 \] \[\cos^2(\frac{A}{2})=\frac{1 }{2} \rightarrow \cos(\frac{A}{2})=\sqrt{\frac{1 }{2} }\] \[\rightarrow \cos(\frac{A}{2})=\frac{1 }{\sqrt 2}\] \[\rightarrow \cos(\frac{A}{2})=\cos45^0 \rightarrow \frac{A}{2} =45^0\] \[A=2 \times 45^0 \rightarrow A = 90^0\]

OpenStudy (anonymous):

@basith

OpenStudy (akashdeepdeb):

As you are not there I'll just write the proof \[\cos 2\theta= \cos(\theta+ \theta)= \cos \theta.\cos \theta - \sin \theta. \sin \theta = \cos^2 \theta - \sin^2 \theta = \cos^2 \theta - (1-\cos^2 \theta)\] \[2\cos^2 \theta - 1 = \cos 2 \theta\] NOW YOU CAN USE THIS! OR You can use @dpasingh 's method.

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

thnx guys i understud.... :)

OpenStudy (akashdeepdeb):

:)